# AuroraCap: Efficient, Performant Video Detailed Captioning and a New Benchmark

### v1

@@ 4 Conclusion
In this paper, we first introduce AuroraCap, a efficient video detailed captioner based on large multimodal model. By leveraging the token merging strategy, we significantly reduce the computational overhead without compromising performance. We also present VDC, a novel video detailed captioning benchmark designed to evaluate comprehensive and coherent textual descriptions of video content. For better evaluating, We propose VDCscore , a new LLM-assisted metric with divide-and-conquer strategy. Our extensive evaluation on various video and image captioning benchmarks demonstrated that AuroraCap achieves competitive results, even outperforming state-of-the-art models in some tasks. We also conduct thorough ablation studies to validate the effectiveness of token merging and other aspects of our model. We found that the current model performs poorly in terms of the trade-off between performance and the scale of input tokens. Additionally, there is still room for improvement in camera handling and detailed captioning. We hope that VDC can bring new insights to the video detailed captioning task.
@@ Acknowledgments
We acknowledge Pika Lab for the computing resources support. We are also thankful to Xtuner  [^1] GitHub Repository: https://github.com/InternLM/xtuner, lmms-eval  [^2] GitHub Repository: https://github.com/EvolvingLMMs-Lab/lmms-eval, and SGLang  [^3] GitHub Repository: https://github.com/sgl-project/sglang for their well-developed and user-friendly codebase. We would like to thank Enxin Song for her contribution to the benchmark processing. We would like to thank Professor Yilun Du, Professor Saining Xie, and Professor Christopher D. Manning for their valuable discussions and revisions of this paper.
@@
We have made several efforts to ensure the reproducibility of our work. All the key implementation details, including the architecture of our model, the training procedures, and hyperparameter settings, are described in supplementary meterial Section E. The introduction of the used evaluation benchmarks and settings are in Section F. The introduction of the used evaluation benchmarks and settings are in Section F. Prompt template of VDC generation is in Section H. Question-answer pairs generation prompt template of VDCscore is in Section I. And calculation details of Elo ranking is in Section L. We have also outlined any hardware configurations and computation requirement used for our experiments in Figure D6 to further support reproducibility.
@@
  • Iiterature review about the related works in Section A.
  • Benchmark comparison between VDC and other video captioning benchmarks in Section B.
  • The calculation and the visualization examples of token merging in Section C.
  • More ablation studies for AuroraCap in Section D.
  • The training settings for AuroraCap in Section E.
  • The introduction of the used evaluation benchmarks and settings in Section F.
  • More limitations of AuroraCap at current stage in Section G.
  • Prompt template of VDC generation in Section H.
  • Question-answer pairs generation prompt template of VDCscore in Section I.
  • More statistics information of VDC in Section J.
  • More statistics information of VDCscore in Section K.
  • Calculation details of Elo ranking in Section L.
  • Case studies among several baselines in Section M.
Appendix A Related Works
@@
The general goal of video captioning is understanding a video and describing it with natural language. Unlike image captioning, video captioning requires the description also on the temporal dimension (e.g., human action, camera and object movement). The current datasets including videos in the domain of human activities Caba Heilbron et al. (2015), cooking scene Zhou et al. (2018); Iashin & Rahtu (2020), movie Song et al. (2023), and open domain Chen & Dolan (2011). In Table 1, we show the summary of current video captioning benchmarks. Most of them are with short caption, which is not suitable for evaluating video detailed captioning task. There are several widely used metrics to evaluate the correctness of generated caption, such as BLEU Papineni et al. (2002), ROUGE-L Lin (2004), CIDEr Vedantam et al. (2015), SPICE Anderson et al. (2016), and BERTScore Zhang et al. (2019). Dense Video Captioning Yang et al. (2023a); Zhou et al. (2024) is also a captioning task but further requires localizing the events temporally. In this paper, we focus on the detailed description of a short video clip, where there are no scene changes or camera switches.
@@
With the develop of LLMs and LMMs Maaz et al. (2023); Zhang et al. (2023b); Li et al. (2023f); Song et al. (2023); Zhang et al. (2023a); Song et al. (2024); Zhao et al. (2024a); Zhou et al. (2024); team (2024); Jin et al. (2024); Beyer et al. (2024); Shang et al. (2024); Cai et al. (2024); Cheng et al. (2024); Li et al. (2024c), many recent works have explored adapting them into video understanding field (e.g., Video-LLaMA Li et al. (2023f), Video-LLaVA Lin et al. (2023a), VideoChat Li et al. (2023b), Vista-LLaMA Ma et al. (2023), LLaVA-Hound Zhang et al. (2024c), Koala Tan et al. (2024), Elysium Wang et al. (2024a), and MovieChat Song et al. (2023; 2024)). Thanks to this flexible design, the models can combine pretrained knowledge with minimal trainable parameters. Instead of requiring large-scale training, using only a small amount of high-quality training data can even achieve better results. Most of the existing models use additional parameters for temporal modeling. There are also some interesting observations that we can actually build advancing LMMs without additional parameters for temporal modeling Xu et al. (2024a); Lin et al. (2023b); Liu et al. (2024a); Zhang et al. (2024a); Ren et al. (2023) or even without further training with video-text data Wu (2024). Recent works (e.g., FreeVA Wu (2024), LLaVA-Next Liu et al. (2024a), VLIA Lin et al. (2023b), and PLLaVA Xu et al. (2024a)) also find that the vanilla LLaVA-like model pretrained on high-quality image instruction data can also be a strong video understanding model. FreeVA further observe that using existing video instruction tuning data like Video-ChatGPT 100K Maaz et al. (2023) to tune LMMs may not necessarily lead to improvements. As the concurrent work, we also observe this phenomenon and proceeded to develop a video detailed captioning baseline training based on the LLaVA architecture.
@@
As for the benchmark , inspired by LLaVA Liu et al. (2024b), Video-ChatGPT Maaz et al. (2023) introduces a 100K video clips with text instructions with the first vLMMs benchmark evaluation system powered by LLMs. MovieChat-1K Song et al. (2023) and CinePile Rawal et al. (2024) are question-answering based benchmark for long-form video understanding. Shot2Story20K Han et al. (2023) comprises videos with 2 to 8 shots each sourced for our dataset from the public video benchmark HDvila100M Xue et al. (2022). However, currently there is no video benchmark available to evaluate video detailed captioning tasks like IIW Garg et al. (2024) did in image captioning field. In this paper, our work fills this gap.
Appendix B Benchmark Comparison
@@
We compare our proposed VDCwith some examples from several video captioning benchmarks as shown in Figure B1, Figure B2, and Figure B3. The corresponding captions are shown as followings:
Figure B1: Video example of MSR-VTT Xu et al. (2016) benchmark.
Figure B2: Video example of VATEX Wang et al. (2019) benchmark.
Figure B3: Video example of VDC benchmark sampled from Ego4D Grauman et al. (2022).
@@
Appendix C Token Merging Visualization
Figure C4: Token merging visualization. From top to bottom, the image IDs are COCO:COCO-train2014-000000247906, VG:2331508, SA-1B:sa-393200, OCR-VQA:71789111. From left to right, the number of tokens representing the images are 490, 154, 18, and 6.
@@ Appendix C Token Merging Visualization
We show token merging visualization examples from COCO Lin et al. (2014), VG Krishna et al. (2017b), SA-1B Kirillov et al. (2023), and OCR-VQA Mishra et al. (2019) in Figure C4, since we only conduct frame-wise operation. Features extracted by CLIP ViT-H model  [^4] Huggingface Model: https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14-378. We conduct image captioning for the examples given final token numbers by 7B AuroraCap as following. We remove line breaks in generated caption for a better reading experience.
@@
    Describe this image in detail. (Figure C4 Row 1)
@@
    What is the number in the picture? (Figure C4 Row 2)
@@
    What is the relative position of the white and red cars in the picture? Which one in the left? (Figure C4 Row 3)
@@ Appendix D Ablation Studies
[…]
@@
Inspired by Slowfast-LLaVA Xu et al. (2024b), we explore whether combining frames with low and high $R_{vtk}$ can enhance performance. In practice, we don’t conduct token merging in the first frame and concatenate them with the merged tokens from subsequent frames. We apply this strategy to both video captioning and video question answering tasks, comparing performance with and without the inclusion of full first-frame visual tokens. As illustrated in Table D1, slowfast inference brings marginal performance improvement in video question answering tasks or even drop in video captioning tasks but with more computing cost. Therefore, by default, we don’t using slowfast inference for video detailed captioning.
We also present the performance curve with and without the inclusion of full first-frame visual tokens as the visual token kept ratio varies during inference across multiple video understanding tasks. As illustrated in Figure D9 and Figure D10, despite the inclusion of full first-frame visual tokens, slowfast inference does not consistently result in significantly positive effects on performance. In some cases, the incorporation of full first-frame visual tokens even worsens performance degradation as the kept ratio decreases, particularly in video captioning tasks.
Table D1: Ablation on slowfast inference for AuroraCap-7B. We present the average performance among different token merging ratio on various video understanding benchmarks. We show that slowfast inference brings marginal performance improvement in video question answering tasks or even drop in video captioning tasks but with more computing cost.
@@
Efficiency.
To assess the inference speed, we utilize the inference time per video question-answering pair in seconds (TPV) as an evaluative metric. SGLang is an accelerated serving framework for LLMs and multimodal LLMs. We consider four settings including with or without token merging and SGLang. Figure D5 indicates the minimum TPV achievable in each settings across seven video understanding datasets. Reducing the visual tokens and using SGLang result in excellent inference times per video question-answering pair while all the datasets with short video and question inputs. In contrast, maintaining full visual tokens or omitting the use of SGLang results in comparatively slower performance, demonstrating the superior inference efficiency of AuroraCap. For each input, we process the video at a resolution of 378 $\times$ 378 and sample 8 frames using single H100 GPU.
Figure D5: Comparison between different inference settings: $\mathbb{A}$: $R_{vtk}=1.0$, without SGLang, $\mathbb{B}$: $R_{vtk}=0.1$, without SGLang, $\mathbb{C}$: $R_{vtk}=1.0$, with SGLang, $\mathbb{D}$: $R_{vtk}=0.1$, with SGLang. The number indicates the maximum inference time in seconds for each benchmark.
@@
  • Setting $\mathbb{B}$: Following Slowfast-LLaVA Xu et al. (2024b), we retain full visual tokens in the first frame and concatenate with merged tokens using $R_{vtk}$ of 0.1.
@@
  • Setting $\mathbb{D}$: Most videos in the training data have no more than 8 key frames, while a subset (mainly from ShareGPT4Video Chen et al. (2024d)) contains significantly more. We first exclude this subset from end-to-end training and then use it to train solely the LLM, enhancing its ability to handle multi-frame inputs.
@@
We implement these training strategies, track training costs in H100 hours, and evaluate across various video understanding tasks. As shown in Figure D6, while training with an $R_{vtk}$ of 1.0 improves performance, it significantly increases training time. Surprisingly, mixing lower and higher visual token ratios during training offers no significant advantage. Training only the LLM under the two settings results in a performance drop, indicating that enhancing long video understanding still requires collaboration with the finetuning the visual encoder. Therefore, we choose Setting $\mathbb{C}$ as the final training strategy.
Figure D6: Comparison between different training strategy in Language stage. We take Accuracy for Question-Answering tasks and CIDEr for captioning tasks as the evaluation metric and present the performance percentage. We choose Setting $\mathbb{C}$ as the final training strategy as shown with *. The number shows the maximum value for each benchmark.
Figure D7: Ablation study of token merging on image captioning on Flickr Young et al. (2014), NoCaps Agrawal et al. (2019), COCO-Cap Lin et al. (2014), visual question answering in GQA Hudson & Manning (2019), MMMU Yue et al. (2023), VQAv2 Goyal et al. (2017). We found that token merging significantly reduces the number of tokens while maintaining minimal performance drop, and even showing improvement in some tasks. We highlight the token merging ratio when achieving 90% and 80% performance with the dash line and filled area.
Figure D8: Ablation study of token merging on video captioning on MSRVTT Xu et al. (2016), VATEX Wang et al. (2019), video question answering on MSRVTT-QA Xu et al. (2016), ActivityNet-QA Yu et al. (2019), MSVD-QA Xu et al. (2017), iVQA Liu et al. (2018). We found that token merging significantly reduces the number of tokens while maintaining minimal performance drop, and even showing improvement in some tasks. We highlight the token merging ratio when achieving 90% and 80% performance with the dash line and filled area.
Figure D9: Comparison between performance with and without the inclusion of full first-frame visual tokens during inference on video captioning on MSRVTT Xu et al. (2016), VATEX Wang et al. (2019), video question answering on MSRVTT-QA Xu et al. (2016), ActivityNet-QA Yu et al. (2019), MSVD-QA Xu et al. (2017), iVQA Liu et al. (2018). As the visual token kept ratio varies, the blue curve indicates performance with slowfast inference, while the red curve represents performance without slowfast inference.
@@
We use the the training loss among the last ten iterations in original LLaVA alignment pretraining stage to guidance the ViT and LLM backbones selection as shown in Table D2.
Table D2: Final training loss during pretraining stage with original LLaVA pretraining data.

### v2

@@ 4 Conclusion
In this paper, we first introduce AuroraCap, a efficient video detailed captioner based on large multimodal model. By leveraging the token merging strategy, we significantly reduce the computational overhead without compromising performance. We also present VDC, a novel video detailed captioning benchmark designed to evaluate comprehensive and coherent textual descriptions of video content. For better evaluating, We propose VDCscore , a new LLM-assisted metric with divide-and-conquer strategy. Our extensive evaluation on various video and image captioning benchmarks demonstrated that AuroraCap achieves competitive results, even outperforming state-of-the-art models in some tasks. We also conduct thorough ablation studies to validate the effectiveness of token merging and other aspects of our model. We found that the current model performs poorly in terms of the trade-off between performance and the scale of input tokens. Additionally, there is still room for improvement in camera handling and detailed captioning.
@@ Acknowledgments
We acknowledge Pika Labs for the computing resources support. We are also thankful to Xtuner***GitHub Repository: https://github.com/InternLM/xtuner, lmms-eval***GitHub Repository: https://github.com/EvolvingLMMs-Lab/lmms-eval, and SGLang***GitHub Repository: https://github.com/sgl-project/sglang for their well-developed and user-friendly codebase. We would like to thank Enxin Song for her contribution to the benchmark processing. We would like to thank Professor Yilun Du, Professor Saining Xie, and Professor Christopher D. Manning for their valuable discussions and revisions of this paper.
@@
We have made several efforts to ensure the reproducibility of our work. All the key implementation details, including the architecture of our model, the training procedures, and hyperparameter settings, are described in supplementary meterial Section 11. The introduction of the used evaluation benchmarks and settings are in Section 12. The introduction of the used evaluation benchmarks and settings are in Section 12. Prompt template of VDC generation is in Section 15. Question-answer pairs generation prompt template of VDCscore is in Section 16. And calculation details of Elo ranking is in Section 19. We have also outlined any hardware configurations and computation requirement used for our experiments in Figure 20 to further support reproducibility. The code for model training, evaluation, deployment, as well as model weight, benchmark, and training dataset of all the training stages will be released with the paper.
@@
  • Iiterature review about the related works in Section 5.
  • More results on long video benchmarks in Section 6.
  • N-gram based evaluation metrics in Section 7.
  • Benchmark comparison between VDC and other video captioning benchmarks in Section 8.
  • The calculation and the visualization examples of token merging in Section 9.
  • More ablation studies for AuroraCap in Section 10.
  • The training settings for AuroraCap in Section 11.
  • The introduction of the used evaluation benchmarks and settings in Section 12.
  • More limitations of AuroraCap at current stage in Section 14.
  • Additional evaluation metrics analysis in Section 13.
  • Prompt template of VDC generation in Section 15.
  • Question-answer pairs generation prompt template of VDCscore in Section 16.
  • More statistics information of VDC in Section 17.
  • More statistics information of VDCscore in Section 18.
  • Calculation details of Elo ranking in Section 19.
  • Case studies among several baselines in Section 20.
  • Prompt template of predicted answers extraction in Section 21.
  • Prompt template of correctness evaluation in Section 22.
5 Related Works
@@
The general goal of video captioning is understanding a video and describing it with natural language. Unlike image captioning, video captioning requires the description also on the temporal dimension (e.g., human action, camera and object movement). The current datasets including videos in the domain of human activities [11], cooking scene [168, 47], movie [114], and open domain [14]. In Table 1, we show the summary of current video captioning benchmarks. Most of them are with short caption, which is not suitable for evaluating video detailed captioning task. There are several widely used metrics to evaluate the correctness of generated caption, such as BLEU [99], ROUGE-L [76], CIDEr [127], SPICE [3], and BERTScore [160]. Dense Video Captioning [147, 169] is also a captioning task but further requires localizing the events temporally. In this paper, we focus on the detailed description of a short video clip, where there are no scene changes or camera switches.
@@
With the develop of LLMs and LMMs [87, 158, 72, 114, 155, 115, 162, 169, 120, 49, 8, 110, 12, 24, 63], many recent works have explored adapting them into video understanding field (e.g., Video-LLaMA [72], Video-LLaVA [75], VideoChat [64], Vista-LLaMA [85], LLaVA-Hound [159], Koala [117], Elysium [129], and MovieChat [114, 115]). Thanks to this flexible design, the models can combine pretrained knowledge with minimal trainable parameters. Instead of requiring large-scale training, using only a small amount of high-quality training data can even achieve better results. Most of the existing models use additional parameters for temporal modeling. There are also some interesting observations that we can actually build advancing LMMs without additional parameters for temporal modeling [141, 77, 81, 156, 106] or even without further training with video-text data [136]. Recent works (e.g., FreeVA [136], LLaVA-Next [81], VLIA [77], and PLLaVA [141]) also find that the vanilla LLaVA-like model pretrained on high-quality image instruction data can also be a strong video understanding model. FreeVA further observe that using existing video instruction tuning data like Video-ChatGPT 100K [87] to tune LMMs may not necessarily lead to improvements. As the concurrent work, we also observe this phenomenon and proceeded to develop a video detailed captioning baseline training based on the LLaVA architecture. Chat-UniVi [49] employ a set of dynamic visual tokens to uniformly represent images and videos. However, for current video inputs, when we sample at 1 FPS, the frames are already sufficiently sparse; frame-to-frame similarity is quite low except in static scenes. On the other hand, if we use keyframes as input, the frame similarity will be even lower. Under these conditions, we believe that strong temporal merging should be avoided, as it is likely to impair the model’s core capabilities, especially since the CLIP vision encoder trained on images.
@@
As for the benchmark, inspired by LLaVA [82], Video-ChatGPT [87] introduces a 100K video clips with text instructions with the first vLMMs benchmark evaluation system powered by LLMs. MovieChat-1K [114] and CinePile [104] are question-answering based benchmark for long-form video understanding. Shot2Story20K [42] comprises videos with 2 to 8 shots each sourced for our dataset from the public video benchmark HDvila100M [143]. However, currently there is no video benchmark available to evaluate video detailed captioning tasks like IIW [38] and other datasets [66] did in image captioning field. Other works [150] focus on the specific domain like Ego. In this paper, our work fills this gap. As for metric, LLaVA-Hound-DPO [159] uses similar divide-and-conquer strategy to build a preference dataset, ensuring that instructional data remains factually consistent with detailed captions. Our goal for VDCscore, however, is to evaluate the accuracy of generated captions through QA pairs generated from ground truth.
6 More Results on Long-form Video Benchmarks
Figure 7: Model performance on MovieChat-1K [114] benchmark.
Model Degign for Long Video Breakpoint Global
Video-LLaMA $\times$ 39.1 51.7
VideoChat $\times$ 46.1 57.8
TimeChat $\checkmark$ 46.1 73.8
VideoChatGPT $\times$ 48.0 47.6
MovieChat $\checkmark$ 48.3 62.3
MovieChat+ $\checkmark$ 49.6 71.2
AuroraCap (ours) $\times$ 52.6 59.7
Figure 8: Model performance on Egoschema [89] benchmark.
Model Design for Long Video Acc.
Random Choice $\times$ 20.0
FrozenBLM $\times$ 26.9
mPLUG-Owl $\times$ 31.1
TimeChat $\times$ 33.0
Video-LLAVA $\checkmark$ 38.4
LLAMA-VID $\checkmark$ 39.5
LLAVA-NEXT-Video $\checkmark$ 43.9
VideoLLAMA2 $\checkmark$ 51.7
MovieChat $\checkmark$ 53.5
Human Performance 76.2
AuroraCap (ours) $\times$ 46.0
Figure 9: Token merging ablation curve on Egoschema [89], with best performance at a visual token kept ratio of 0.6 to 0.7.
Figure 10: Correlation between n-gram evaluation scores and human elo ranking on the VDC.
We take MovieChat-1K [114], Egoschema [88] as the representative long video benchmark, and compare with existing models. As shown in the following Table 8 and Table 8, AuroraCap achieves comparable performance even without training on long videos and special design. We also plot the visual token kept ratio curve on EgoSchema as Figure 10, demonstrating the stability and effectiveness of our token merging method for longer video sequences.
7 N-gram Based Evaluation Results
A variant of VDCscore could be developed to assess final triplets (<question, correct answer, predicted answer>) by employing a n-gram based evaluation approach as shown in Table 9 with correlation shown in Figure 10.
Table 9: Evaluation results of various models on VDC based on ROUGE for answer matching.
Model Avg. Camera Short Background Main Object Detailed
LLaMA-VID 6.66 9.53 5.12 5.30 6.07 7.29
Video-ChatGPT-7B 6.85 9.20 5.37 6.40 6.29 6.97
ViLA-7B 6.68 7.88 5.16 6.29 6.35 7.74
Video-LLAVA-7B 6.79 8.67 5.25 6.09 6.64 7.28
LLAVA-1.5-7B 6.34 8.31 5.40 5.80 5.52 6.65
LongVA-7B 6.70 8.37 4.99 6.11 6.79 7.23
LLAVA-1.5-13B 6.34 8.64 5.55 5.57 5.15 6.79
LLAVA-NeXT-V7B 6.69 8.61 5.30 6.35 6.27 6.93
LLAVA-1.6-7B 6.45 8.08 5.38 5.79 5.82 7.20
LLAVA-1.6-13B 6.18 7.31 5.28 6.00 5.29 7.01
ShareGPT4Video-8B 6.78 8.46 5.73 5.82 6.54 7.36
LLAVA-OV-7B 6.59 8.09 5.66 5.86 5.58 7.75
InternVL-2-8B 6.82 8.14 5.63 5.95 6.53 7.86
AuroraCap-7B 7.60 9.55 5.66 6.79 7.42 8.58
Gemini-1.5 Pro 7.59 8.78 6.25 7.29 7.70 7.94
8 Benchmark Comparison
@@
We compare our proposed VDCwith some examples from several video captioning benchmarks as shown in Figure 11, Figure 12, and Figure 13. The corresponding captions are shown as followings:
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Video example of MSR-VTT [140] benchmark.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Video example of VATEX [133] benchmark.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Video example of VDC benchmark sampled from Ego4D [40].
@@
9 Token Merging Visualization
Figure 14: Token merging visualization. From top to bottom, the image IDs are COCO:COCO-train2014-000000247906, VG:2331508, SA-1B:sa-393200, OCR-VQA:71789111. From left to right, the number of tokens representing the images are 490, 154, 18, and 6.
@@ 9 Token Merging Visualization
We show token merging visualization examples from COCO [78], VG [57], SA-1B [54], and OCR-VQA [94] in Figure 14, since we only conduct frame-wise operation. Features extracted by CLIP ViT-H model ***Huggingface Model: https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14-378. We conduct image captioning for the examples given final token numbers by 7B AuroraCap as following. We remove line breaks in generated caption for a better reading experience.
@@
    Describe this image in detail. (Figure 14 Row 1)
@@
    What is the number in the picture? (Figure 14 Row 2)
@@
    What is the relative position of the white and red cars in the picture? Which one in the left? (Figure 14 Row 3)
@@ 10 Ablation Studies
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 15: Ablation study of token merging on image captioning on Flickr [151], NoCaps [1], COCO-Cap [78], visual question answering in GQA [45], MMMU [153], VQAv2 [39]. We found that token merging significantly reduces the number of tokens while maintaining minimal performance drop, and even showing improvement in some tasks. We highlight the token merging ratio when achieving 90% and 80% performance with the dash line and filled area.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 16: Ablation study of token merging on video captioning on MSRVTT [140], VATEX [133], video question answering on MSRVTT-QA [140], ActivityNet-QA [152], MSVD-QA [138], iVQA [79]. We found that token merging significantly reduces the number of tokens while maintaining minimal performance drop, and even showing improvement in some tasks. We highlight the token merging ratio when achieving 90% and 80% performance with the dash line and filled area.
@@
As a core strategy of AuroraCap, token merging plays a significant role in reducing the number of visual tokens. We conduct extensive ablation studies to explore the impact of the token kept ratio $R_{vtk}$ in terms of performance across multiple tasks including image captioning, visual question answering, video captioning, and video question answering as shown in Figure 15 and Figure 16. We define the performance percentage as the proportion between the highest and lowest values on the entire performance curve. We identify the minimum retention thresholds for achieving 90% and 80% performance. As shown in Figure 15, while the performance of AuroraCap generally declines with fewer visual tokens across most benchmarks, it remains relatively stable at higher retention levels. Most models maintain satisfactory performance ($>$ 80%) even with only 0.4 of $R_{vtk}$, highlighting the efficiency of our approach. Visual token retention thresholds vary by task complexity, with more visually demanding tasks needing higher retention of visual tokens. For instance, CIDEr [127] on COCO-Cap [128] maintains over 90% performance with an $R_{vtk}$ of 0.3, whereas accuracy on GQA [45] drops to 90% when the $R_{vtk}$ is reduced to 0.8. Unlike image understanding, the optimal performance across most video understanding benchmarks occurs at a relatively low $R_{vtk}$ as depicted in Figure 16. And for MSR-VTT [140], VATEX [133], and ActivityNet-QA [152], even achieve better results at extremely low $R_{vtk}$ ($<$ 0.1). It indicates that comparing to image, video input have higher redundancy. Note that AuroraCap focuses on spatial visual token merging, while the temporal features introduce additional complexity to explore the token merging laws. Appendix 9 shows more calculation details and the visualization results of token merging.
Efficiency.
To assess the inference speed, we utilize the inference time per video question-answering pair in seconds (TPV) as an evaluative metric. SGLang is an accelerated serving framework for LLMs and multimodal LLMs. We consider four settings including with or without token merging and SGLang. Figure 17 indicates the minimum TPV achievable in each settings across seven video understanding datasets. Reducing the visual tokens and using SGLang result in excellent inference times per video question-answering pair while all the datasets with short video and question inputs. In contrast, maintaining full visual tokens or omitting the use of SGLang results in comparatively slower performance, demonstrating the superior inference efficiency of AuroraCap. For each input, we process the video at a resolution of 378 $\times$ 378 and sample 8 frames using single H100 GPU.
Figure 17: Comparison between different inference settings: $\mathbb{A}$: $R_{vtk}=1.0$, without SGLang, $\mathbb{B}$: $R_{vtk}=0.1$, without SGLang, $\mathbb{C}$: $R_{vtk}=1.0$, with SGLang, $\mathbb{D}$: $R_{vtk}=0.1$, with SGLang. The number indicates the maximum inference time in seconds for each benchmark.
@@
Inspired by Slowfast-LLaVA [142], we explore whether combining frames with low and high $R_{vtk}$ can enhance performance. In practice, we don’t conduct token merging in the first frame and concatenate them with the merged tokens from subsequent frames. We apply this strategy to both video captioning and video question answering tasks, comparing performance with and without the inclusion of full first-frame visual tokens. As illustrated in Table 10, slowfast inference brings marginal performance improvement in video question answering tasks or even drop in video captioning tasks but with more computing cost. Therefore, by default, we don’t using slowfast inference for video detailed captioning.
Table 10: Ablation on slowfast inference for AuroraCap-7B. We present the average performance among different token merging ratio on various video understanding benchmarks. We show that slowfast inference brings marginal performance improvement in video question answering tasks or even drop in video captioning tasks but with more computing cost.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 18: Comparison between performance with and without the inclusion of full first-frame visual tokens during inference on VDC. As the visual token kept ratio varies, the blue curve indicates performance with slowfast inference, while the red curve represents performance without slowfast inference.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 19: Comparison between performance with and without the inclusion of full first-frame visual tokens during inference on video captioning on MSRVTT [140], VATEX [133], video question answering on MSRVTT-QA [140], ActivityNet-QA [152], MSVD-QA [138], iVQA [79]. As the visual token kept ratio varies, the blue curve indicates performance with slowfast inference, while the red curve represents performance without slowfast inference.
We also present the performance curve with and without the inclusion of full first-frame visual tokens as the visual token kept ratio varies during inference across multiple video understanding tasks. As illustrated in Figure 19 and Figure 18, despite the inclusion of full first-frame visual tokens, slowfast inference does not consistently result in significantly positive effects on performance. In some cases, the incorporation of full first-frame visual tokens even worsens performance degradation as the kept ratio decreases, particularly in video captioning tasks.
@@
  • Setting $\mathbb{B}$: Following Slowfast-LLaVA [142], we retain full visual tokens in the first frame and concatenate with merged tokens using $R_{vtk}$ of 0.1.
@@
  • Setting $\mathbb{D}$: Most videos in the training data have no more than 8 key frames, while a subset (mainly from ShareGPT4Video [19]) contains significantly more. We first exclude this subset from end-to-end training and then use it to train solely the LLM, enhancing its ability to handle multi-frame inputs.
@@
We implement these training strategies, track training costs in H100 hours, and evaluate across various video understanding tasks. As shown in Figure 20, while training with an $R_{vtk}$ of 1.0 improves performance, it significantly increases training time. Surprisingly, mixing lower and higher visual token ratios during training offers no significant advantage. Training only the LLM under the two settings results in a performance drop, indicating that enhancing long video understanding still requires collaboration with the finetuning the visual encoder. Therefore, we choose Setting $\mathbb{C}$ as the final training strategy.
Figure 20: Comparison between different training strategy in Language stage. We take Accuracy for Question-Answering tasks and CIDEr for captioning tasks as the evaluation metric and present the performance percentage. We choose Setting $\mathbb{C}$ as the final training strategy as shown with *. The number shows the maximum value for each benchmark.
@@
We use the the training loss among the last ten iterations in original LLaVA alignment pretraining stage to guidance the ViT and LLM backbones selection as shown in Table 11.
Table 11: Final training loss during pretraining stage with original LLaVA pretraining data.